# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v2

@@ 1 Introduction
Regularized estimators are a fundamental technique for addressing overfitting in machine learning and data science, particularly in modern high-dimensional datasets; thus, understanding their theoretical behavior has become all the more important. To this end, the prototypical problem of ridge regression has gained renewed attention, with multiple works characterizing its out-of-sample risk in high dimensions, c.f., [7, 18]. Additionally, consistent estimates for this out-of-sample risk, crucial for optimizing the regularization parameter $\lambda$, have been established in the form of leave-one-out cross-validation (LOOCV) and generalized cross-validation (GCV) [45, 29, 3].
@@
Crucially, all aforementioned analyses assume that samples are independent and identically distributed (i.i.d.) draws from some distribution with sufficient regularity, such as light tails. These assumptions often fail in practice, such as in financial returns, neuroscience, and climate data [14, 48, 37]. In such cases, cross-validation techniques can become biased (Theorem 1, Remark 3), necessitating alternatives. Only recently, [4] considered independent but non-identically distributed samples and [2] explored Gaussian designs with both sample and feature dependence. Understanding cross-validation in high dimensions under more general sample dependencies and for designs with heavier tails than sub-Gaussians remains a significant challenge. This paper takes the first step in addressing this issue.
We assess the impact of sample dependence and heavy tails in ridge regression and cross-validation using an alternative random matrix ensemble for the covariate distribution. Instead of assuming a Gaussian matrix or i.i.d. rows from a fixed distribution,
@@
we require the singular value decomposition of the design $\mathbf{X}$ to be right-rotationally invariant (see Definition 1). We characterize the ridge regression risk for these designs and introduce ROTI-GCV, a new GCV-inspired framework for consistently estimating the out-of-sample risk. Simulations show that ROTI-GCV outperforms traditional GCV and LOOCV in situations with correlated and heavy-tailed observations.
Right-rotationally invariant designs facilitate tractable analysis while capturing structured forms of sample dependencies and heavy-tailed covariates.
These designs have gained significant attention as alternatives to i.i.d. designs for theoretical analysis in numerous studies
@@
[38, 26, 31, 9, 12, 39, 25, 22, 10, 46]. Recent work [8, 41] showed that, under mild conditions, properties of convex estimators across various designs are well-approximated by those under a right-rotationally invariant design with the same spectrum, indicating a broad universality class for these distributions.
@@
Formally, we operate in the proportional asymptotics regime, where the dimension $p$ scales proportionally with the sample size $n$. This regime, rooted in probability theory and statistical physics, has received significant attention due to its ability to accurately capture empirical phenomena [18, 35, 36, 24]. Furthermore, results derived under this regime require minimal assumptions on the signal structure, resulting in theory and methods with broad practical utility [34]. In regards to cross-validation for i.i.d. samples, [18, 30, 40] established that classical $k$-fold cross-validation is inconsistent estimates for the out-of-sample error in this high-dimensional regime, whereas LOOCV and GCV remain consistent, prompting our study of GCV. However, our work focuses on non-i.i.d. samples and non-Gaussian tailed designs, areas not sufficiently addressed in prior work.
@@ 2.1 Right-rotationally invariant designs
Some examples of right-rotationally invariant designs are as follows (note that an i.i.d. entries standard Gaussian design is a member of this class, but the class includes many non-i.i.d. and non-Gaussian designs as well; see Appendix B.3 for proofs of invariance):
  1. (i)
@@
    Autocorrelated data: The rows of $\mathbf{X}$ can be drawn from an autocorrelated series, where $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$.
  2. (ii)
@@
    $t$-distributed data: The rows of $\mathbf{X}$ can be drawn from a multivariate $t$ distribution with as few as $3$ degrees of freedom, capturing heavy-tailed data such as financial returns [14].
  3. (iii)
@@
    Products of Gaussian matrices: $\mathbf{X}=\mathbf{X}_{1}\mathbf{X}_{2}\cdot\cdots\cdot\mathbf{X}_{k}$, where $\mathbf{X}_{1}$ has $n$ rows and $\mathbf{X}_{k}$ has $p$ columns, while the remaining dimensions are arbitrary. See [16, 17] for connections to linear neural networks.
  4. (iv)
@@
    Matrix-Normals: $\mathbf{X}\sim\mathsf{MN}(\mathbf{0},\boldsymbol{\Sigma}^{\mathrm{row}},\boldsymbol{\Sigma}^{\mathrm{col}})$, where $\boldsymbol{\Sigma}^{\mathrm{row}}$ can be arbitrary while $\boldsymbol{\Sigma}^{\mathrm{col}}\sim\mathsf{InvWishart}(\mathbf{I}_{p},(1+\delta)p)$, for any $\delta>0$. See Appendix B.5 for details on distribution notation.
  5. (v)
@@
    Equicorrelated data: $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$.
  6. (vi)
@@
    Spiked matrices: $\mathbf{X}=\lambda\mathbf{V}\mathbf{W}^{\top}+\mathbf{G}$, where $\mathbf{V}\in\mathbb{R}^{n\times r}$ and $\mathbf{W}\in\mathbb{R}^{p\times r}$ are the first $r$ columns of two Haar matrices, and $G_{ij}\sim\mathsf{N}(0,1)$.
@@
See [23, Figure 1] for more such examples. Setting (ii) has heavier tails than supported by existing GCV theory; settings (i, iii, v, vi) have sample dependence; setting (iv) has both.
Right-rotationally invariant ensembles allow us enormous flexibility—they can simultaneously capture dependent rows and heavy-tails by allowing a rather general class of singular value distributions.
@@
The spectrum of right-rotationally invariant designs can essentially be arbitrary, and allowing for significant generalization from prior i.i.d. designs used to study the behavior of LOOCV and GCV in high dimensions. Furthermore, the universality class of these designs is extremely broad, as established in [8, 41].
To simplify some statements, we introduce some notation. Let $m_{\mathbf{D}}(z)=\frac{1}{p}\sum_{i=1}^{p}\frac{1}{D_{ii}^{2}-z}$ and $v_{\mathbf{D}}(z)=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{D_{ii}^{2}-z}$, with $D_{ii}=0$ if $i>\min(n,p)$.
@@
Also of use will be the derivatives $m^{\prime}_{\mathbf{D}}$ and $v^{\prime}_{\mathbf{D}}$. Note $m_{\mathbf{D}}$ is the Stieltjes transform of the empirical measure of $(\mathbf{D}^{\top}\mathbf{D}_{ii})_{i=1}^{p}$; $v_{\mathbf{D}}$ is the companion Stieltjes transform.
@@ 2.2 Problem Setup
with $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$, and $\boldsymbol{\epsilon},\boldsymbol{\beta}\in\mathbb{R}^{p}$. Hence $n$ is the sample size, $p$ is both the data dimension and the parameter count, $\boldsymbol{\beta}$ is the signal, and $\boldsymbol{\epsilon}$ is a component-wise independent noise vector, where each entry has mean 00 and variance $\sigma^{2}$. We study the expected out-of-sample risk on an independent draw from a new, potentially different right-rotationally invariant design, $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $\widetilde{\mathbf{X}}\in\mathbb{R}^{n^{\prime}\times p}$. This corresponds to the setting of learning a model on one interdependent population, and using the model on a separate interdependent population. We analyze the conditional risk
@@
Such conditional risks have been examined in prior works [18, 29].
We focus on the risk of the ridge estimator
@@
    We consider a sequence of right-rotationally invariant designs $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n},\dots$ where each $\mathbf{X}_{i}\sim\mathbf{Q}_{i}^{\top}\mathbf{D}_{i}\mathbf{O}_{i}$, and a sequence of signal vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\dots,\boldsymbol{\beta}_{n},\dots$ such that for all $n$, $\boldsymbol{\beta}_{n}\in\mathbb{R}^{p}$ and $\|\boldsymbol{\beta}_{n}\|=r\sqrt{n}$; $\mathbf{y}_{i}$ is generated by Eq. 1. In the sequel, we drop the dependence on $n$ and write $\mathbf{X},\mathbf{y}$ and $\boldsymbol{\beta}$ with the indexing implicit. We refer to $(\mathbf{X},\mathbf{y})$ as the training data.
@@
    lim $\limsup\lambda_{\max}(\mathbf{D}_{n})<C$ a.s. for some constant $C$.
@@
    $\widetilde{\mathbf{X}}$ is an independent draw from a potentially different right-rotationally invariant ensemble, meaning $\widetilde{\mathbf{X}}\sim\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $(\widetilde{\mathbf{Q}},\widetilde{\mathbf{D}},\widetilde{\mathbf{O}})$ independent of $\mathbf{X}$. $\widetilde{\mathbf{Q}}$ and $\widetilde{\mathbf{D}}$ need not be distributed the same as $\mathbf{Q},\mathbf{D}$; $\widetilde{\mathbf{O}}$ must be Haar.
@@
To clarify why the risk in Eq. 2 is relevant, if one samples $(\widetilde{\mathbf{X}},\widetilde{\mathbf{y}})$ according to Eq. (1), then $R_{\mathbf{X},\mathbf{y}}(\lambda)=\mathbb{E}[\|\widetilde{\mathbf{y}}-\widetilde{\mathbf{X}}\hat{\boldsymbol{\beta}}_{\lambda}\|^{2}\mid\mathbf{X},\mathbf{y}]-\sigma^{2}$, meaning the risk defined corresponds, up to a factor of $\sigma^{2}$, to the mean-squared error on the test set.
@@ 3 Generalized Cross-Validation
Prior work [45, 29] established that for many designs with i.i.d. samples, LOOCV and GCV estimate the out-of-sample error consistently under the proportional asymptotics regime, while $k$-fold CV does not. In our setting, LOOCV is intractable without additional assumptions on $\mathbf{Q}$ (see discussion in Appendix B.6). However, GCV is tractable, motivating our study here. Initially introduced as a technique for selecting the regularization parameter in smoothing splines [6], GCV’s optimality has been shown in various tasks [21, 15]. It is also a consistent tuning technique for kernel ridge regression [47, 43, 44].
The GCV cross-validation metric for ridge regression is as follows [13]:
@@
$\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})=r^{2}\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)+\sigma^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))$ (6)
@@
In Theorem 2, one can interpret the term involving $r^{2}$ as the bias of the estimator, and the term with $\sigma^{2}$ as its variance. As such, one can then interpret Theorem 1 as showing how GCV reconstructs the bias and variance from the data. See Appendix A.4 for explicit derivations of the bias and variance terms.
@@
It is shown in [29] that for i.i.d. designs, $\Delta_{n}(\lambda):=\mathsf{GCV}_{n}(\lambda)-R_{\mathbf{X},\mathbf{y}}(\lambda)-\sigma^{2}$ approaches zero uniformly over compact intervals [^1] The extra $\sigma^{2}$ term is discussed in Remark 1.. Their proof of this relies explicitly on identities of the Stieltjes transform given by the Silverstein equation [33]. These identities do not exist for right-rotationally invariant designs, and thus the GCV becomes biased for these designs, meaning $\Delta_{n}(\lambda)\not\to 0$. The more the spectrum departs from that of the i.i.d. setting, the more biased the estimates.

### v3

@@ 1 INTRODUCTION
Regularized estimators are fundamental for modern high-dimensional statistics. In this context, the prototypical problem of ridge regression has gained renewed attention, with multiple works characterizing its out-of-sample risk in high dimensions, c.f., Dobriban and Wager, (2018); Hastie et al., (2022). Additionally, consistent estimates for this out-of-sample risk, crucial for optimizing the regularization parameter $\lambda$, have been established in the form of leave-one-out cross-validation (LOOCV) and generalized cross-validation (GCV) (Xu et al., , 2021; Patil et al., , 2021; Atanasov et al., 2024b, ).
@@
Crucially, all aforementioned analyses assume that samples are independent and identically distributed (i.i.d.) draws from some distribution with sufficient regularity, such as light tails. These assumptions often fail in practice, such as in financial returns, neuroscience, and climate data (Grobys, , 2021; Zscheischler et al., , 2021; Tagliazucchi et al., , 2013). In such cases, cross-validation techniques can become biased (Theorem 1, Remark 3), necessitating alternatives. Only recently, Bigot et al., (2024) considered independent but non-identically distributed samples and Atanasov et al., 2024a explored Gaussian designs with both sample and feature dependence. Understanding cross-validation in high dimensions under more general sample dependencies and for designs with heavier tails than sub-Gaussians remains a significant challenge. This paper takes the first step in addressing this issue.
@@
We assess the impact of sample dependence and heavy tails in ridge regression and cross-validation using an alternative random matrix ensemble for the covariate distribution. Instead of assuming a Gaussian matrix or i.i.d. rows from a fixed distribution, we require the singular value decomposition of the design $\mathbf{X}$ to be right-rotationally invariant (see Definition 1). We characterize the ridge regression risk for these designs and introduce ROTI-GCV, a new GCV-inspired framework for consistently estimating the out-of-sample risk. Simulations show that ROTI-GCV outperforms traditional GCV and LOOCV in situations with correlated and heavy-tailed observations.
@@
Right-rotationally invariant designs facilitate tractable analysis while capturing structured forms of sample dependencies and heavy-tailed covariates. These designs have gained significant attention as alternatives to i.i.d. designs for theoretical analysis in numerous studies (Takeda et al., , 2006; Ma and Ping, , 2017; Rangan et al., , 2019; Fan, , 2022; Gerbelot et al., , 2020; Takeuchi, , 2019; Liu et al., , 2022; Li et al., , 2023; Fan and Wu, , 2021; Xu et al., , 2022). Recent works (Dudeja et al., , 2023; Wang et al., , 2023) showed that, under mild conditions, properties of convex estimators across various designs are well-approximated by those under a right-rotationally invariant design with the same spectrum, indicating a broad universality class for these distributions.
@@
Formally, we operate in the proportional asymptotics regime, where the dimension $p$ scales proportionally with the sample size $n$. This regime, rooted in probability theory and statistical physics, has received significant attention due to its ability to accurately capture empirical phenomena (Hastie et al., , 2022; Sur and Candès, , 2019; Sur et al., , 2019; Liang and Sur, , 2022). Furthermore, results derived under this regime require minimal assumptions on the signal structure, resulting in theory and methods with broad practical utility Song et al., (2024). In regards to cross-validation for i.i.d. samples, Hastie et al., (2022); Rahnama Rad and Maleki, (2018); Wang et al., (2018) established that classical $k$-fold cross-validation inconsistently estimates the out-of-sample error in this high-dimensional regime, whereas LOOCV and GCV remain consistent, prompting our study of GCV. However, our work focuses on non-i.i.d. samples and non-Gaussian tailed designs, areas not sufficiently addressed in prior work.
@@ 2.1 Right-rotationally invariant designs
Some examples of right-rotationally invariant designs are as follows (note that an i.i.d. entries standard Gaussian design is a member of this class, but the class includes many non-i.i.d. and non-Gaussian designs as well; see Appendix 8.3 for proofs of invariance):
@@
(i) Autocorrelated data: The rows of $\mathbf{X}$ can be drawn from an autocorrelated series, where $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\mathbf{z}_{i}$, with $\mathbf{z}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$.
@@
(ii) $t$-distributed data: The rows of $\mathbf{X}$ can be drawn from a multivariate $t$ distribution with as few as $3$ degrees of freedom, capturing heavy-tailed data such as financial returns (Grobys, , 2021).
@@
(iii) Products of Gaussian matrices: $\mathbf{X}=\mathbf{X}_{1}\mathbf{X}_{2}\cdot\cdots\cdot\mathbf{X}_{k}$, where $\mathbf{X}_{1}$ has $n$ rows and $\mathbf{X}_{k}$ has $p$ columns, while the remaining dimensions are arbitrary. See Hanin and Nica, (2020); Hanin and Paouris, (2021) for connections to linear neural networks.
@@
(iv) Matrix-Normals: $\mathbf{X}\sim\mathsf{MN}(\mathbf{0},\boldsymbol{\Sigma}^{\mathrm{row}},\boldsymbol{\Sigma}^{\mathrm{col}})$, where $\boldsymbol{\Sigma}^{\mathrm{row}}$ can be arbitrary while $\boldsymbol{\Sigma}^{\mathrm{col}}\sim\mathsf{InvWishart}(\mathbf{I}_{p},(1+\delta)p)$, for any $\delta>0$. See Appendix 8.5 for details on distribution notation.
@@
(v) Equicorrelated data: $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$.
@@
(vi) Spiked matrices: $\mathbf{X}=\lambda\mathbf{V}\mathbf{W}^{\top}+\mathbf{G}$, where $\mathbf{V}\in\mathbb{R}^{n\times r}$ and $\mathbf{W}\in\mathbb{R}^{p\times r}$ are the first $r$ columns of two Haar matrices, and $G_{ij}\sim\mathsf{N}(0,1)$.
@@
See (Li and Sur, , 2023, Figure 1) for more such examples. Setting (ii) has heavier tails than supported by existing GCV theory; settings (i, iii, v, vi) have sample dependence; setting (iv) has both.
@@
Right-rotationally invariant ensembles allow us enormous flexibility—they can simultaneously capture dependent rows and heavy-tails by allowing a rather general class of singular value distributions. The spectrum of right-rotationally invariant designs can essentially be arbitrary, and allowing for significant generalization from prior i.i.d. designs used to study the behavior of LOOCV and GCV in high dimensions. Furthermore, the universality class of these designs is extremely broad, as established in Dudeja et al., (2023); Wang et al., (2023).
@@
To simplify some statements, we introduce some notation. Let $m_{\mathbf{D}}(z)=\frac{1}{p}\sum_{i=1}^{p}\frac{1}{D_{ii}^{2}-z}$ and $v_{\mathbf{D}}(z)=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{D_{ii}^{2}-z}$, with $D_{ii}=0$ if $i>\min(n,p)$. Also of use will be the derivatives $m^{\prime}_{\mathbf{D}}$ and $v^{\prime}_{\mathbf{D}}$. Note $m_{\mathbf{D}}$ is the Stieltjes transform of the empirical measure of $(\mathbf{D}^{\top}\mathbf{D}_{ii})_{i=1}^{p}$; $v_{\mathbf{D}}$ is the Stieltjes transform of $(\mathbf{D}\mathbf{D}^{\top}_{ii})_{i=1}^{n}$.
@@ 2.2 Problem Setup
with $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$, and $\boldsymbol{\epsilon},\boldsymbol{\beta}\in\mathbb{R}^{p}$. Hence $n$ is the sample size, $p$ is both the data dimension and the parameter count, $\boldsymbol{\beta}$ is the signal, and $\boldsymbol{\epsilon}$ is a component-wise independent noise vector, where each entry has mean 0 and variance $\sigma^{2}$. We study the expected out-of-sample risk on an independent draw from a new, potentially different right-rotationally invariant design, $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $\widetilde{\mathbf{X}}\in\mathbb{R}^{n^{\prime}\times p}$. This corresponds to learning a model on one interdependent population, and using it on a separate interdependent population. We analyze the conditional risk
@@
Such conditional risks have been examined in prior works (Hastie et al., , 2022; Patil et al., , 2021). We focus on the risk of the ridge estimator
@@
    We consider a sequence of right-rotationally invariant designs $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n},\dots$ where each $\mathbf{X}_{i}\sim\mathbf{Q}_{i}^{\top}\mathbf{D}_{i}\mathbf{O}_{i}$, and a sequence of signal vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\dots,\boldsymbol{\beta}_{n},\dots$ such that for all $n$, $\boldsymbol{\beta}_{n}\in\mathbb{R}^{p}$ and $\|\boldsymbol{\beta}_{n}\|=r\sqrt{n}$; $\mathbf{y}_{i}$ is generated by Eq. (1). In the sequel, we drop the dependence on $n$ and write $\mathbf{X},\mathbf{y}$ and $\boldsymbol{\beta}$ with the indexing implicit. We refer to $(\mathbf{X},\mathbf{y})$ as the training data.
@@
    $\limsup\lambda_{\max}(\mathbf{D}_{n})<C$ a.s. for some constant $C$.
@@
    $\widetilde{\mathbf{X}}$ is an independent draw from a potentially different right-rotationally invariant ensemble.
@@
[…]
@@ 3 GENERALIZED CROSS-VALIDATION
Prior work (Xu et al., , 2021; Patil et al., , 2021) established that for many designs with i.i.d. samples, LOOCV and GCV estimate the out-of-sample error consistently in the proportional asymptotics regime, while $k$-fold CV does not. In our setting, LOOCV is intractable without additional assumptions on $\mathbf{Q}$ (see Appendix 8.6). However, GCV is tractable, motivating our study here. Initially introduced as a technique for selecting the regularization parameter in smoothing splines Craven and Wahba, (1978), GCV’s optimality and consistency has been shown in various tasks (Li, , 1986; Gu and Ma, , 2005; Zhang et al., , 2015; Xu et al., , 2018, 2019). The GCV cross-validation metric for ridge regression is as follows (Golub et al., , 1979):
@@
$\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})=r^{2}\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)\\ +\sigma^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))$ (6)
@@
In Theorem 2, one can interpret the term involving $r^{2}$ as the bias of the estimator, and the term with $\sigma^{2}$ as its variance. As such, one can then interpret Theorem 1 as showing how GCV reconstructs the bias and variance from the data.
@@
It is shown in Patil et al., (2021) that for i.i.d. designs, $\Delta_{n}(\lambda):=\mathsf{GCV}_{n}(\lambda)-R_{\mathbf{X},\mathbf{y}}(\lambda)-\sigma^{2}$ approaches zero uniformly over compact intervals [^1] The extra $\sigma^{2}$ term is discussed in Remark 1.. Their proof of this relies explicitly on identities of the Stieltjes transform given by the Silverstein equation (Silverstein, , 1995). These identities do not exist for right-rotationally invariant designs, and thus the GCV becomes biased for these designs, meaning $\Delta_{n}(\lambda)\not\to 0$. The more the spectrum departs from that of the i.i.d. setting, the more biased the estimates.